OpenStudy (pokemon23):

Solve for b: A=1/2h(b+c) show all work

OpenStudy (saifoo.khan):

Mini will give it a try.

OpenStudy (pokemon23):

your smartest guy I know ;)

OpenStudy (pokemon23):

alos the guy with the bicycle pic

OpenStudy (saifoo.khan):

lol, dont forget sat. he is the smartest here.

OpenStudy (pokemon23):

I know the guy with bicycle

OpenStudy (anonymous):

is this \[ a=\frac{1}{2}h(b+c)\]?

OpenStudy (pokemon23):

that the correct question :0

OpenStudy (anonymous):

multiply by 2, divide by h, subtract c

OpenStudy (pokemon23):

multiply two by 1/2

OpenStudy (pokemon23):

I got 1

OpenStudy (mimi_x3):

b = c -2a/h ?

OpenStudy (saifoo.khan):

\[A=\frac 12h(b+c)\] \[2A = h(b+c)\] divide h on both sides, \[\frac{2A}{h} = b+c\] \[b = \frac{2A}{h} - c\]

OpenStudy (mimi_x3):

yay ! i got it right ^_^

OpenStudy (pokemon23):

that nice mimi

OpenStudy (saifoo.khan):

lol.

OpenStudy (saifoo.khan):

My answer was incorrect @mini.

OpenStudy (mimi_x3):

what ? incorrect >

OpenStudy (saifoo.khan):

Answer.